# Embodied AI: Emerging Risks and Opportunities for Policy Action

### v1

@@ 3.2 What are the most significant gaps?
First, there are significant gaps concerning robust certification processes for different EAI modalities. Regulating AVs is straightforward in many senses due to their defined operational domains (e.g. cars usually stay on roads). The EU’s MR mandates certification of safety for a broader array of EAI modalities (with carve-outs for AVs or aerial drones) and stipulates that EAI “shall not…perform actions beyond its defined task and movement space.” Current drone regulations address many aspects of aerial drones, but this regulation largely focuses on drones lacking true autonomous navigation capabilities. Future instantiations of EAI, however, will likely have significantly expanded freedom of movement, enabling them to enter private residences and commercial establishments and conduct surveillance in schools or public areas. These expanded domains require thoughtful processes to certify the operational safety of EAIs—such frameworks do not yet exist. Expecting existing consumer-safety laboratories, which currently test the safety of machine components like materials and locking mechanisms, to evaluate the safety of EAI is unrealistic. Basic questions such as identifying the relevant regulator are a key starting point—to what extent should there be EAI-specific consumer protection boards with AI expertise, or should existing third-party testing laboratories take on this responsibility?
@@
Secondly, once a suitable apparatus is in place, EAI capabilities should be measured with reliable and valid evaluations and benchmarks. To date, few of these benchmarks exist Sermanet et al. (2025), despite the existence of a range of benchmarks for virtual AI systems noa (2024e). Voluntary standards can provide technical guidance, but the lack of laws enforcing benchmarking and evaluations across all four key risk areas (physical, informational, economic, and social) is a critical policy gap. Evaluations could cover a range of considerations outlined in our taxonomy, for example evaluating the robustness of simulation-to-real protocols (the “reality gap”), the conformity of EAIs with their stated operational domain, robustness to jailbreaking, cybersecurity measures, alignment between software and hardware capabilities, and hardware durability and reliability, among other areas.
@@
Thirdly, policies or frameworks currently devoted to post-deployment EAI monitoring are unclear or lacking in detail. These sorts of oversight and monitoring mechanisms have been highlighted for other AI systems Dunlop and Stein (2024). However, current regulations requiring EAIs to include “black boxes” that record and preserve data in the event of accidents, crashes, or misuse are hazy. The EU’s MR mandates that data about safety-related decision-making processes is kept for a year after its collection noa (2023). At the same time, the EU’s AI Act provides contradictory guidance that high-risk AI systems must retain this information for at least six months noa (2024d). These types of recording systems, in addition to live data monitoring, can enhance system safety and aid post-incident investigations Kukkala et al. (2022); Mathew et al. (2025). The EU’s MR also states that “it shall be possible at all times to correct the machinery…to maintain its inherent safety,” but this notion of oversight requires significant clarification for highly- or fully-autonomous systems. Does this involve the ability to tweak EAI actions in real-time, send new model updates over the air, or some other intervention? Who is conducting this monitoring—users, the government, or a private, delegated oversight entity (perhaps AI-driven itself)?
@@
There are likewise stark gaps in policies addressing economic and social risks from EAI. Although EAI systems could cause mass labor displacement, proposals to distribute economic benefits are still in their infancy Yelizarova (2025); Merola (2022). EAI could lead to many positive economic and social outcomes as well, but policymakers must then ensure that economies around the world have the sovereign systems (data centers, energy production, EAI hardware) necessary to seize EAI’s benefits. No well-articulated policy on this, whether national or regional, yet exists. Similarly, policies addressing social issues related to trust and human-EAI attachment are currently scant Bao et al. (2023). More broadly, there are significant policy gaps at the intersection of EAI and AGI. For example, should an EAI be allowed to build other EAIs? Should a country developing AGI in embodied form automatically and freely share the technology with the rest of the world? There has likewise been little policy attention devoted to EAI defensive acceleration. If AGI is as powerful as some observers imagine, it seems possible or even plausible that AGI could help solve a raft of open governance questions and issues, particularly in the defense arena Clifford (2024). Yet what does it mean for a society to be protected by an army of EAIs? Policymakers must urgently consider how EAI should be developed, deployed, and integrated into societal structures to address the broad array of currently neglected challenges mentioned here.
@@ 4.1 Invest in EAI safety research
Based on the risk taxonomy described in Section 2, we recommend significant and increased research be devoted to EAI safety. For example, robotics and machine-learning researchers can further efforts to make hardware actuators less susceptible to hacking and malfunction through physical design and formal methods Basin (2021). Building benchmarks and evaluations of EAI capabilities and behavior is a particularly promising area of EAI safety research. Most AI benchmarks and evaluations today specifically target the virtual aspects of AI Grey and Segerie (2025), although recent progress has been made in EAI-specific research, such as at Google DeepMind Sermanet et al. (2025). Researchers should and build on this progress by developing EAI evaluations and benchmarks that span a broad set of tasks and task types Yin et al. (2025); Choi et al. (2024), similar to the work being done by the RoboArena team Atreya et al. (2025). For example, specific attention could focus on ensuring the safety of EAIs acting in multi-agent systems or swarms. EAIs will almost certainly navigate complex environments with other EAIs that do not share the same goals, so inter-agent collaboration and coordination are paramount to avoid poor outcomes, as recently highlighted by the World Economic Forum and researchers from Cooperative AI Larsen et al. (2024); Hammond et al. (2025). Beyond physical risks, benchmarks and evaluations should also address issues related to privacy and cybersecurity, for example by building upon zero-knowledge proof research from other AI domains Peng et al. (2025). We also need benchmarks that stress-test the joint vision–language–action loop—measuring, for instance, whether a VLA model’s visual prompt leads to safe, context-aware behavior across edge cases. Benchmarks and assessments will not address every risk raised above—particularly socioeconomic considerations—but they are a critical step towards minimizing many risks from EAI.
@@ 4.2 Create robust certification requirements before EAI deployment
National bodies should mandate that EAI systems pass safety evaluations and are certified for public use. EAIs should have clear ‘model cards’ describing how an EAI system was trained (e.g. what sorts of data were used, how a model performs on safety benchmarks), in which domains it was designed to operate, and what safety measures the manufacturer has taken to ensure its safe operation. If desired, policymakers could then mandate that EAIs be limited to legally operating within the specified domains (as in the EU’s MR), potentially aided by remote identification requirements similar to those currently applicable to drones. This model card approach would borrow from the frontier safety frameworks that many leading AI labs have implemented Buhl et al. (2025). This regime could be enforced via audits of EAI manufacturers and developers Mökander (2023). Policymakers should clarify which entity is responsible for verifying and validating EAI safety, possibly by establishing a national laboratory for EAI testing as part of existing AI Safety Institutes or by assigning this responsibility to private-sector actors. For a stricter approach, policymakers could also decide that only EAI systems quantifiably proven safe, as currently being investigated in conjunction with the UK’s ARIA funding agency, can be deployed Dalrymple (2024). Policymakers should also ensure that this certification regime incorporates different categories of requirements based on potential risk, as risk from EAI depends heavily on the sensitivity of an EAI’s deployed context and its capabilities, as noted in the recent RAND report “Averting a Robot Catastrophe.” Vermeer et al. (2025) For example, certifying EAI safety for an EAI version of a children’s toy and a autonomous limousine should involve different safety testing requirements and thresholds. Designing the exact certification regimen and constructing these different categories will require significant collaboration with technical experts. Policymakers should ensure that this sort of regulation is reasonably limited in scope and supports beneficial innovation while mitigating risks.
@@ 4.3 Promote industry-led standards to address EAI risks
Industrial and standards bodies can push forward EAI safety efforts in tandem with legislative approaches. These standards bodies can develop robust updates to existing standards and create dynamic new standards that address the increased capabilities of EAI. Existing standards are grounded in today’s robotic capabilities. Still, even recent updates fail to incorporate or address how highly or entirely autonomous robots capable of advanced reasoning will affect industrial and service applications. Notably, in May 2025 the ISO announced intentions to create a new standard for humanoid robots, which should encompass a range of additional form factors, autonomy, and use cases noa (2025h). These standards should address a range of technical protocols (e.g. cybersecurity and jailbreak-proof), while also mandating that EAI systems are equipped with tools to foster transparency and accountability. For instance, future standards should mandate that EAI systems are equipped with “black boxes” that record the EAI system’s sensor input and, if possible, the reasoning in the minutes preceding an adverse event. These black boxes will raise privacy concerns of their own, as they would create an extra attack surface for data exfiltration. The tension between accountability and privacy should be acknowledged and addressed in future discussions.
@@ 4.5 Plan and prepare for the transformative economic and social effects of EAI
Similarly, policymakers must better prepare for transformative social effects Kulveit et al. (2025). Given the capital-intensive nature of EAI systems, it is plausible that EAI power and access could be concentrated in the hands of a select few Ford (2016). Policymakers should draft options to combat this social power concentration, perhaps through targeted taxation mechanisms Thuemmel (2023); Ahn (2024); Mazur (2019). Policymakers should also fund research on how to mitigate adverse emotional dependencies between EAIs and humans. EAI deployment is ultimately (for now) a human decision, so national and international policymakers should consider whether some domains should be entirely off-limits for EAI interaction. Organizations such as the OECD, GPAI, or the nascent UN AI Panel and Dialogue should prioritize action-oriented EAI dialogue on these pressing social issues, as EAI will impact people worldwide, not just in today’s robotics hotspots.
@@ 5 Discussion and Limitations
Skeptics could argue that many issues discussed in this paper are just inherent LLM risks. In other words, if we make LLMs safe, we will solve EAI risks. We acknowledge that there is considerable overlap between risks from AI and EAI Weidinger et al. (2022). However, EAI raises specific concerns about safety at the application and organizational level, in addition to the fundamental model level Mökander et al. (2024). Many risks are modified, magnified, or made more urgent by the presence of EAIs in the physical world. EAI creates immediate risks (e.g. a malfunction during surgery could result in a patient’s death), pervasive (e.g. with continuous presence in households, workplaces, and public areas), and growing with scale (e.g. swarms of public-safety EAI systems coordinating to enforce martial law). These risks will likely be exacerbated as EAI capabilities, autonomy, and deployment increase. For example, an EAI could impair a baby’s healthy development Sharkey and Sharkey (2012) or be hacked and cause a deadly crash in ways inapplicable to purely virtual AIs Yousseef et al. (2024). Improving LLM safety alone is helpful but would fail to cover these—and many other—critical EAI scenarios.
@@ 6 Conclusion
We have argued that policymakers should encourage the development of effective benchmarks, evaluations, and safety protocols for the responsible deployment of EAI; ensure safety certification for a range of EAI form factors, capabilities, and operational domains; reevaluate liability paradigms; confront labor displacement; and address larger societal issues, such as human-EAI attachment. These recommendations aim to provide sensible first steps that can guide and encourage safe EAI innovation with minimal downside. Many other risks necessitate minor tweaks or adjustments to existing policies—for example, preventing privacy violations from EAIs in public will require thoughtful integration with existing laws such as the MR and GDPR. Zooming out, research on EAI safety in general should be significantly expanded, and policymakers must collaborate with robotics and machine-learning researchers and practitioners to translate findings from academic and industry research into policy. EAI is rapidly advancing, and its risks are quickly becoming real; policymakers must urgently address policy gaps to mitigate these critical risks.

### v2

@@ 3.2 What are the most significant gaps?
First, there are significant gaps concerning robust certification processes for different EAI modalities. Regulating AVs is straightforward in many senses due to their defined operational domains (e.g. cars usually stay on roads). The EU’s MR mandates certification of safety for a broader array of EAI modalities (with carve-outs for AVs or aerial drones) and stipulates that EAI “shall not…perform actions beyond its defined task and movement space.” Current drone regulations address many aspects of aerial drones, but this regulation largely focuses on drones lacking true autonomous navigation capabilities. Future instantiations of EAI, however, will likely have significantly expanded freedom of movement, enabling them to enter private residences and commercial establishments and conduct surveillance in schools or public areas. These expanded domains require thoughtful processes to certify the operational safety of EAI systems—such frameworks do not yet exist. Expecting existing consumer-safety laboratories, which currently test the safety of machine components like materials and locking mechanisms, to evaluate the safety of EAI sytems is unrealistic. Basic questions such as identifying the relevant regulator are a key starting point—to what extent should there be EAI-specific consumer protection boards with AI expertise, or should existing third-party testing laboratories take on this responsibility?
@@
Secondly, once a suitable apparatus is in place, EAI capabilities should be measured with reliable and valid evaluations and benchmarks. To date, few of these benchmarks exist Sermanet et al. (2025), despite the existence of a range of benchmarks for virtual AI systems noa (2024e). Voluntary standards can provide technical guidance, but the lack of laws enforcing benchmarking and evaluations across all four key risk areas (physical, informational, economic, and social) is a critical policy gap. Evaluations could cover a range of considerations outlined in our taxonomy, for example evaluating the robustness of simulation-to-real protocols (the “reality gap”), the conformity of EAI systems with their stated operational domain, robustness to jailbreaking, cybersecurity measures, alignment between software and hardware capabilities, and hardware durability and reliability, among other areas.
@@
Thirdly, policies or frameworks currently devoted to post-deployment EAI monitoring are unclear or lacking in detail. These sorts of oversight and monitoring mechanisms have been highlighted for other AI systems Dunlop and Stein (2024). However, current regulations requiring EAI systems to include “black boxes” that record and preserve data in the event of accidents, crashes, or misuse are hazy. The EU’s MR mandates that data about safety-related decision-making processes is kept for a year after its collection noa (2023). At the same time, the EU’s AI Act provides contradictory guidance that high-risk AI systems must retain this information for at least six months noa (2024d). These types of recording systems, in addition to live data monitoring, can enhance system safety and aid post-incident investigations Kukkala et al. (2022); Mathew et al. (2025). The EU’s MR also states that “it shall be possible at all times to correct the machinery…to maintain its inherent safety,” but this notion of oversight requires significant clarification for highly- or fully-autonomous systems. Does this involve the ability to tweak EAI actions in real-time, send new model updates over the air, or some other intervention? Who is conducting this monitoring—users, the government, or a private, delegated oversight entity (perhaps AI-driven itself)?
@@
There are likewise stark gaps in policies addressing economic and social risks from EAI. Although EAI could cause mass labor displacement, proposals to distribute economic benefits are still in their infancy Yelizarova (2025); Merola (2022). EAI could lead to many positive economic and social outcomes as well, but policymakers must then ensure that economies around the world have the sovereign systems (data centers, energy production, EAI hardware) necessary to seize EAI’s benefits. No well-articulated policy on this, whether national or regional, yet exists. Similarly, policies addressing social issues related to trust and human-EAI attachment are currently scant Bao et al. (2023). More broadly, there are significant policy gaps at the intersection of EAI and AGI. For example, should an EAI system be allowed to build other EAI systems? Should a country developing AGI in embodied form automatically and freely share the technology with the rest of the world? There has likewise been little policy attention devoted to EAI defensive acceleration. If AGI is as powerful as some observers imagine, it seems possible or even plausible that AGI could help solve a raft of open governance questions and issues, particularly in the defense arena Clifford (2024). Yet what does it mean for a society to be protected by an army of EAI systems? Policymakers must urgently consider how EAI should be developed, deployed, and integrated into societal structures to address the broad array of currently neglected challenges mentioned here.
@@ 4.1 Invest in EAI safety research
Based on the risk taxonomy described in Section 2, we recommend significant and increased research be devoted to EAI safety. For example, robotics and machine-learning researchers can further efforts to make hardware actuators less susceptible to hacking and malfunction through physical design and formal methods Basin (2021). Building benchmarks and evaluations of EAI capabilities and behavior is a particularly promising area of EAI safety research. Most AI benchmarks and evaluations today specifically target the virtual aspects of AI Grey and Segerie (2025), although recent progress has been made in EAI-specific research, such as at Google DeepMind Sermanet et al. (2025). Researchers should and build on this progress by developing EAI evaluations and benchmarks that span a broad set of tasks and task types Yin et al. (2025); Choi et al. (2024), similar to the work being done by the RoboArena team Atreya et al. (2025). For example, specific attention could focus on ensuring the safety of EAI systems acting in multi-agent systems or swarms. EAI systems will almost certainly navigate complex environments with others that do not share the same goals, so inter-agent collaboration and coordination are paramount to avoid poor outcomes, as recently highlighted by the World Economic Forum and researchers from Cooperative AI Larsen et al. (2024); Hammond et al. (2025). Beyond physical risks, benchmarks and evaluations should also address issues related to privacy and cybersecurity, for example by building upon zero-knowledge proof research from other AI domains Peng et al. (2025). We also need benchmarks that stress-test the joint vision–language–action loop—measuring, for instance, whether a VLA model’s visual prompt leads to safe, context-aware behavior across edge cases. Benchmarks and assessments will not address every risk raised above—particularly socioeconomic considerations—but they are a critical step towards minimizing many risks from EAI.
@@ 4.2 Create robust certification requirements before EAI deployment
National bodies should mandate that EAI systems pass safety evaluations and are certified for public use. EAI systems should have clear ‘model cards’ describing how they were trained (e.g. what sorts of data were used, how a model performs on safety benchmarks), in which domains it was designed to operate, and what safety measures the manufacturer has taken to ensure its safe operation. If desired, policymakers could then mandate that EAI systems be limited to legally operating within the specified domains (as in the EU’s MR), potentially aided by remote identification requirements similar to those currently applicable to drones. This model card approach would borrow from the frontier safety frameworks that many leading AI labs have implemented Buhl et al. (2025). This regime could be enforced via audits of EAI manufacturers and developers Mökander (2023). Policymakers should clarify which entity is responsible for verifying and validating EAI safety, possibly by establishing a national laboratory for EAI testing as part of existing AI Safety Institutes or by assigning this responsibility to private-sector actors. For a stricter approach, policymakers could also decide that only EAI systems quantifiably proven safe, as currently being investigated in conjunction with the UK’s ARIA funding agency, can be deployed Dalrymple (2024). Policymakers should also ensure that this certification regime incorporates different categories of requirements based on potential risk, as risk from EAI depends heavily on the sensitivity of an EAI’s deployed context and its capabilities, as noted in the recent RAND report “Averting a Robot Catastrophe.” Vermeer et al. (2025) For example, certifying EAI safety for an EAI version of a children’s toy and a autonomous limousine should involve different safety testing requirements and thresholds. Designing the exact certification regimen and constructing these different categories will require significant collaboration with technical experts. Policymakers should ensure that this sort of regulation is reasonably limited in scope and supports beneficial innovation while mitigating risks.
@@ 4.3 Promote industry-led standards to address EAI risks
Industrial and standards bodies can push forward EAI safety efforts in tandem with legislative approaches. These standards bodies can develop robust updates to existing standards and create dynamic new standards that address the increased capabilities of EAI. Existing standards are grounded in today’s robotic capabilities. Still, even recent updates fail to incorporate or address how highly or entirely autonomous robots capable of advanced reasoning will affect industrial and service applications. Notably, in May 2025 the ISO announced intentions to create a new standard for humanoid robots, which should encompass a range of additional form factors, autonomy, and use cases noa (2025h). These standards should address a range of technical protocols (e.g. cybersecurity and jailbreak-proof), while also mandating that EAI systems are equipped with tools to foster transparency and accountability. For instance, future standards should mandate that EAI systems are equipped with “black boxes” that record the system’s sensor input and, if possible, its reasoning in the minutes preceding an adverse event. These black boxes will raise privacy concerns of their own, as they would create an extra attack surface for data exfiltration. The tension between accountability and privacy should be acknowledged and addressed in future discussions.
@@ 4.5 Plan and prepare for the transformative economic and social effects of EAI
Similarly, policymakers must better prepare for transformative social effects Kulveit et al. (2025). Given the capital-intensive nature of EAI systems, it is plausible that EAI power and access could be concentrated in the hands of a select few Ford (2016). Policymakers should draft options to combat this social power concentration, perhaps through targeted taxation mechanisms Thuemmel (2023); Ahn (2024); Mazur (2019). Policymakers should also fund research on how to mitigate adverse emotional dependencies between EAI systems and humans. EAI deployment is ultimately (for now) a human decision, so national and international policymakers should consider whether some domains should be entirely off-limits for EAI interaction. Organizations such as the OECD, GPAI, or the nascent UN AI Panel and Dialogue should prioritize action-oriented EAI dialogue on these pressing social issues, as EAI will impact people worldwide, not just in today’s robotics hotspots.
@@ 5 Discussion and Limitations
Skeptics could argue that many issues discussed in this paper are just inherent LLM risks. In other words, if we make LLMs safe, we will solve EAI risks. We acknowledge that there is considerable overlap between risks from AI and EAI Weidinger et al. (2022). However, EAI raises specific concerns about safety at the application and organizational level, in addition to the fundamental model level Mökander et al. (2024). Many risks are modified, magnified, or made more urgent by its presence in the physical world. EAI creates immediate risks (e.g. a malfunction during surgery could result in a patient’s death), pervasive (e.g. with continuous presence in households, workplaces, and public areas), and growing with scale (e.g. swarms of public-safety EAI systems coordinating to enforce martial law). These risks will likely be exacerbated as EAI capabilities, autonomy, and deployment increase. For example, an EAI could impair a baby’s healthy development Sharkey and Sharkey (2012) or be hacked and cause a deadly crash in ways inapplicable to purely virtual AIs Yousseef et al. (2024). Improving LLM safety alone is helpful but would fail to cover these—and many other—critical EAI scenarios.
@@ 6 Conclusion
We have argued that policymakers should encourage the development of effective benchmarks, evaluations, and safety protocols for the responsible deployment of EAI; ensure safety certification for a range of EAI form factors, capabilities, and operational domains; reevaluate liability paradigms; confront labor displacement; and address larger societal issues, such as human-EAI attachment. These recommendations aim to provide sensible first steps that can guide and encourage safe EAI innovation with minimal downside. Many other risks necessitate minor tweaks or adjustments to existing policies—for example, preventing privacy violations from EAI systems in public will require thoughtful integration with existing laws such as the MR and GDPR. Zooming out, research on EAI safety in general should be significantly expanded, and policymakers must collaborate with robotics and machine-learning researchers and practitioners to translate findings from academic and industry research into policy. EAI is rapidly advancing, and its risks are quickly becoming real; policymakers must urgently address policy gaps to mitigate these critical risks.